How do you convert 300nM (nanomolar) to molarity? And to a %?

Maybe these conversions won't work, but I have to convert this number into two different units.
Answer #1


1. 300 nM = 300*10^-9 M

          = 3*10^-7 M

(nano means 10^-9 )

answer: 3*10^-7 M

2. molarity to w/v%

Molarity = (weight/Mwt)*(1000/V)

Molarity = (w/Mwt)*(100*10/V)

Molarity = (w/v)*(100*10/Mwt)

Molarity(M) = (w/v%)*10/Mwt (this is the relation between molarity(M) and w/v%)

(w/v%) = weight by volume percentage.

Mwt = molecular weight of solute